Find the domain of the function. g(x)= 5/8-5x
Choose the correct domain below.
a) {xIx is a real number and x ≥ 8/5}
b) {xIx is a real number and x ≠ 0}
c) {xIx is a real number and x ≠ 8/5}
b) {xIx is a real number and x ≠ 5}

Answer by stanbon(75887)

Find the domain of the function. g(x)= 5/(8-5x)
The denominator cannot zero.
Solve 8-5x = 0
5x = 8
x = 8/5
-------------------------------------------------
Ans: c
--------
Cheers,
Stan H.
=================================================
Choose the correct domain below.
a) {xIx is a real number and x ≥ 8/5}
b) {xIx is a real number and x ≠ 0}
c) {xIx is a real number and x ≠ 8/5}
b) {xIx is a real number and x ≠ 5}